# Faster Gradient Methods for Highly-smooth Stochastic Bilevel Optimization

### v1

@@ Abstract
This paper studies the complexity of finding an $\epsilon$-stationary point for stochastic bilevel optimization when the upper-level problem is nonconvex and the lower-level problem is strongly convex. Recent work proposed the first-order method, F2SA, achieving the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper complexity bound for first-order smooth problems. This is slower than the optimal $\Omega(\epsilon^{-4})$ complexity lower bound in its single-level counterpart. In this work, we show that faster rates are achievable for higher-order smooth problems. We first reformulate F2SA as approximating the hyper-gradient with a forward difference. Based on this observation, we propose a class of methods F2SA-$p$ that uses $p$th-order finite difference for hyper-gradient approximation and improves the upper bound to $\tilde{\mathcal{O}}(p\epsilon^{4-p/2})$ for $p$th-order smooth problems. Finally, we demonstrate that the $\Omega(\epsilon^{-4})$ lower bound also holds for stochastic bilevel problems when the high-order smoothness holds for the lower-level variable, indicating that the upper bound of F2SA-$p$ is nearly optimal in the highly smooth region $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$.
{NoHyper}
@@ 1 Introduction
Many machine learning problems, such as meta-learning [39], hyper-parameter tuning [5, 16, 36], and adversarial training [18] can be abstracted as solving the following bilevel optimization problem:
@@
We call $f$ and $g$ the upper-level and lower-level functions, respectively, and call $\varphi$ the hyper-objective. In this paper, we consider the most common nonconvex-strongly-convex setting where $f:{\mathbb{R}}^{d_{x}}\rightarrow{\mathbb{R}}$ is smooth and possibly nonconvex, and $g:{\mathbb{R}}^{d_{y}}\rightarrow{\mathbb{R}}$ is smooth jointly in $({\bm{x}},{\bm{y}})$ and strongly convex in $\bm{y}$}. Under the lower-level strong convexity assumption, the implicit function theorem indicates the following closed form of the hyper-gradient [17]:
@@
Following the works in nonconvex optimization [6, 7, 3], we consider the task of finding an $\epsilon$-stationary point of $\varphi$, i.e., a point ${\bm{x}}\in{\mathbb{R}}^{d_{x}}$ such that $\|\nabla\varphi({\bm{x}})\|\leq~\epsilon$. Motivated by many real machine learning tasks, we study the stochastic setting, where the algorithms only have access to the stochastic derivative estimators of both $f$ and $g$.
The first efficient algorithm BSA Ghadimi and Wang [17] for solving the stochastic bilevel problem leverages both stochastic gradient and Hessian-vector-product (HVP) oracles to find an $\epsilon$-stationary point of $\varphi({\bm{x}})$. Subsequently, Ji et al. [21] proposed stocBiO by incorporating multiple enhanced designs to improve the complexity. Both BSA and stocBiO require the stochastic Hessian assumption (5) on the lower-level function, which means $g$ has an unbiased stochastic Hessian estimator with bounded variance. For finite-sum problems, such an assumption is stronger than standard SGD assumptions and equivalent to the mean-squared-smoothness assumption (6) on the lower-level gradient estimator $G$ [2, Observation 1 and 2].
To avoid estimating HVP oracles, Kwon et al. [28] proposed the first fully first-order method F2SA that works under standard SGD assumptions on both $f$ and $g$ (Assumption 2.1). The main idea is to solve the following penalty problem [33, 34, 41]:
@@
where $\lambda$ is taken to be sufficiently large such that $\lambda=\Omega(\epsilon^{-1})$. Thanks to Danskin’s theorem, the gradient of the penalty function in Eq. (3) only involves gradient information. Therefore, F2SA does not require the stochastic Hessian assumptions (5). More importantly, by directly leveraging gradient oracles instead of more expensive HVP oracles, the F2SA is more efficient in practice [40, 42, 22] and it is also the only method that can be scaled to 3232B sized large language model (LLM) training [38].
Kwon et al. [28] proved that the F2SA method finds an $\epsilon$-stationary point of $\varphi({\bm{x}})$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls in the deterministic case and $\tilde{\mathcal{O}}(\epsilon^{-7})$ stochastic first-order oracle (SFO) calls in the stochastic case. Recently, Chen et al. [10] showed the two-time-scale stepsize strategy improves the upper complexity bound of F2SA method to $\tilde{\mathcal{O}}(\epsilon^{-2})$ in the deterministic case, which is optimal up to logarithmic factors. However, the direct extension of their method in the stochastic case leads to the $\tilde{\mathcal{O}}(\epsilon^{-6})$ SFO complexity [10, 29] , which still has a significant gap between the $\Omega(\epsilon^{-4})$ lower bound for SGD [3]. It remains open whether optimal rates for stochastic bilevel problems can be achieved for fully first-order methods.
@@
We can show that F2SA-2 returns an $\epsilon$-estimation to $\nabla\varphi({\bm{x}})$ under the setting $\lambda=\Omega(\epsilon^{-1/2})$ instead of $\Omega(\epsilon^{-1})$ in F2SA, which further improves the SFO complexity of F2SA from $\tilde{\mathcal{O}}(\epsilon^{-6})$ to $\tilde{\mathcal{O}}(\epsilon^{-5})$ for second-order smooth problems. Our idea is generalizable for any $p$th-order problems. We recall that in numerical analysis there exists the $p$th-order central difference that uses $p$ points to construct an estimator to the derivative of a unitary function with $p$th-order error guarantee (Lemma 3.1). Motivated by this fact, we propose the F2SA-$p$ algorithm and show that it allows $\lambda=\Omega(\epsilon^{-1/p})$ for $p$th-order smooth problems, which further leads to the improved $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ SFO complexity for finding an $\epsilon$-stationary point stated by our Theorem 3.1.
To examine the tightness of our upper bounds, we further extend the $\Omega(\epsilon^{-4})$ lower bound for SGD [3] from single-level optimization to bilevel optimization. Note that existing constructions for bilevel lower bound [13, 29] do not satisfy all our smoothness conditions in Definition 2.2. We demonstrate in Theorem 4.1 that a fully separable construction for upper- and lower-level variables can immediately yield a valid $\Omega(\epsilon^{-4})$ lower bound for the problem class we study, showing that F2SA-$p$ is optimal up to logarithmic factors when $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$ (see Remark 3.5). We summarize our main results, including both the lower and upper bounds, in Table 1 and discuss open problems in the following.
@@
Our upper bounds improve known results for high-order smooth problems, but our result still has a gap between the lower bound for $p={\mathcal{O}}(\log\epsilon^{-1}/\log\log\epsilon^{-1})$. Recently, Kwon et al. [29] obtained some preliminary results towards closing this gap for $p=1$, where they showed an $\Omega(\epsilon^{-6})$ lower bound holds under a more adversarial oracle. But it is still open whether their lower bounds can be extended to standard stochastic oracles as they conjectured. Another open problem is the tightness of the condition number dependency shown in Table 1.
@@
F2SA 1st-order [28] $\tilde{\mathcal{O}}({\rm poly}(\kappa)\epsilon^{-7})$
F2SA 1st-order [29] $\tilde{\mathcal{O}}({\rm poly}(\kappa)\epsilon^{-6})$
F2SA 1st-order [10] $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$
F2SA-pp 1st-order Theorem 3.1 $\tilde{\mathcal{O}}(p\kappa^{9+2/p}\epsilon^{-4-2/p})$
Lower Bound $p$th-order in $\bm{y}$} Theorem 4.1 $\Omega(\epsilon^{-4})$
@@ 2 Preliminaries
The goal of bilevel optimization is to minimize the hyper-objective $\varphi({\bm{x}})$, which is in general nonconvex. Since finding a global minimizer of a general nonconvex function requires exponential complexity in the worst case [37, § 1.6], we follow the literature [6, 7] to consider the task of finding an approximate stationary point.
@@
Let $\varphi:{\mathbb{R}}^{d_{x}}\rightarrow{\mathbb{R}}$ be the hyper-objective defined in Eq. (1). We say ${\bm{x}}\in{\mathbb{R}}^{d_{x}}$ is an $\epsilon$-hyper-stationary point if $\|\nabla\varphi({\bm{x}})\|\leq\epsilon$.
Next, we introduce the assumptions used in this paper, which ensure the tractability of the above hyper-stationarity. Compared to [28, 10], we additionally assume the high-order smoothness in lower-level variable $\bm{y}$} to achieve further acceleration.
@@ 2.1 Problem Setup
First of all, we follow the standard assumptions on SGD [3] to assume that the stochastic gradient estimators satisfy the following assumption.
@@
Third, we assume the lower-level function $g({\bm{x}},{\bm{y}})$ is strongly convex in $\bm{y}$}. It guarantees the uniqueness of ${\bm{y}}^{*}({\bm{x}})$ and the tractability of the bilevel problem. Although not all the problems in applications satisfy the lower-level strong convexity assumption, it is impossible to derive dimension-free upper bounds when the lower-level problem is only convex [9, Theorem 3.2]. Hence, we follow most existing works to consider strongly convex lower-level problems.
@@
Fourth, we require the following smoothness assumptions following [17]. According to Eq. (2), these conditions are necessary and sufficient to guarantee the Lipschitz continuity of $\nabla\varphi({\bm{x}})$, which further ensure the tractability of an approximate stationary point of the nonconvex hyper-objective $\varphi({\bm{x}})$ [48, 27].
@@
We refer to the problem class that jointly satisfies all the above Assumption 2.1, 2.2, 2.3 and 2.4 as first-order smooth bilevel problems, for which [29, 10] showed the F2SA method achieves the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper complexity bound. In this work, we show an improved bound under the following additional higher-order smoothness assumption on lower-level variable $\bm{y}$}.
Assumption 2.5 (High order smoothness in $y$).
@@ 2.2 Comparison to Previous Works
All our above assumptions align with [10] except for the additional Assumption 2.5. Since we are not the first work to demonstrate that additional assumptions can lead to acceleration in bilevel optimization, we first give a detailed discussion on other assumptions made in related works to see our differences before we show our improved upper bound.
@@
Ghadimi and Wang [17], Ji et al. [21] assumes the access to a stochastic Hessian estimator ${\bm{H}}({\bm{x}},{\bm{y}})$ such that
@@
Under this assumption, in conjunction with Assumption 2.2, 2.3, and 2.4, Ghadimi and Wang [17] proposed the BSA method and showed that it finds an $\epsilon$-stationary point of $\varphi({\bm{x}})$ with $\tilde{\mathcal{O}}(\epsilon^{-6})$ stochastic gradient oracles and $\tilde{\mathcal{O}}(\epsilon^{-4})$ stochastic HVP oracles. Later, Ji et al. [21] proposed the stocBiO method that requires $\tilde{\mathcal{O}}(\epsilon^{-4})$ stochastic gradient and HVP oracles. Compared to them, we consider the setting where the algorithms only have access to stochastic gradient estimators, and we make no assumptions on the stochastic Hessians.
@@
Besides Assumption 2.1, 2.2, 2.3, 2.4 and the stochastic Hessian assumption (5), Khanduri et al. [25], Yang et al. [46, 47] further assumes that the stochastic estimators to gradients and Hessians are mean-squared smooth:
@@
Under this additional assumption, they proposed faster stochastic methods with upper complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-3})$ via variance reduction [15, 12]. However, variance reduction are typically ineffective in practice [14] since the mean-squared smoothness constants $\bar{L}_{1}$ and $\bar{L}_{2}$ can be arbitrarily worse than the smoothness constants $L_{1}$ and $L_{2}$. In this paper, we only consider the setting without mean-squared smoothness assumptions and study a different acceleration mechanism from variance reduction.
@@
Huang et al. [20] introduced a second-order smoothness assumption similar to but stronger than Assumption 2.5 when $p=2$. Specifically, they assumed the second-order smoothness jointly in $({\bm{x}},{\bm{y}})$ instead of $\bm{y}$} only:
@@
The jointly second-order smoothness (7) ensures that the hyper-objective $\varphi({\bm{x}})$ has Lipschitz continuous Hessians, which further allows the application of known techniques in minimizing second-order smooth objectives. Huang et al. [20] applied the technique from [23, 24, 44, 1] to show that an HVP-based method can find a second-order stationary point in $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity under the deterministic setting, and in $\tilde{\mathcal{O}}(\epsilon^{-4})$ under the stochastic Hessian assumption (5). Yang et al. [45] applied the technique from [31] to accelerate the complexity HVP-based method to $\tilde{\mathcal{O}}({\epsilon}^{-1.75})$ in the deterministic setting. Chen et al. [10] also proposed a fully first-order method to achieve the same $\tilde{\mathcal{O}}({\epsilon}^{-1.75})$ complexity. Compared to these works, our work demonstrates a unique acceleration mechanism in stochastic bilevel optimization that only comes from the high-order smoothness in $\bm{y}$}.

### v2

@@ Abstract
This paper studies the complexity of finding an $\epsilon$-stationary point for stochastic bilevel optimization when the upper-level problem is nonconvex and the lower-level problem is strongly convex. Recent work proposed the first-order method, F2SA, achieving the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper complexity bound for first-order smooth problems. This is slower than the optimal $\Omega(\epsilon^{-4})$ complexity lower bound in its single-level counterpart. In this work, we show that faster rates are achievable for higher-order smooth problems. We first reformulate F2SA as approximating the hyper-gradient with a forward difference. Based on this observation, we propose a class of methods F2SA-$p$ that uses $p$th-order finite difference for hyper-gradient approximation and improves the upper bound to $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ for $p$th-order smooth problems. Finally, we demonstrate that the $\Omega(\epsilon^{-4})$ lower bound also holds for stochastic bilevel problems when the high-order smoothness holds for the lower-level variable, indicating that the upper bound of F2SA-$p$ is nearly optimal in the region $p=\Omega(\log\epsilon^{-1}/\log\log{\epsilon}^{-1})$.
[…]
@@ 1 Introduction
Many machine learning problems, such as meta-learning (Rajeswaran et al., 2019), hyper-parameter tuning (Bao et al., 2021; Franceschi et al., 2018; Mackay et al., 2019), adversarial training (Goodfellow et al., 2020), and reinforcement learning (Yang et al., 2019; Hong et al., 2023; Zeng et al., 2024; Zeng and Doan, 2024) can be abstracted as solving the following bilevel optimization problem:
@@
We call $f$ and $g$ the upper-level and lower-level functions, respectively, and call $\varphi$ the hyper-objective. In this paper, we consider the most common nonconvex-strongly-convex setting where $f:{\mathbb{R}}^{d_{x}}\rightarrow{\mathbb{R}}$ is smooth and possibly nonconvex, and $g:{\mathbb{R}}^{d_{y}}\rightarrow{\mathbb{R}}$ is smooth jointly in $({\bm{x}},{\bm{y}})$ and strongly convex in $\bm{y}$}. Under the lower-level strong convexity assumption, the implicit function theorem (Dontchev and Rockafellar, 2009) indicates the following closed form of the hyper-gradient
@@
Following the works in nonconvex optimization (Carmon et al., 2020; 2021; Arjevani et al., 2023), we consider the task of finding an $\epsilon$-stationary point of $\varphi$, i.e., a point ${\bm{x}}\in{\mathbb{R}}^{d_{x}}$ such that $\|\nabla\varphi({\bm{x}})\|\leq~\epsilon$. Motivated by many real machine learning tasks, we study the stochastic setting, where the algorithms only have access to the stochastic derivative estimators of both $f$ and $g$.
The first efficient algorithm BSA Ghadimi and Wang (2018) for solving the stochastic bilevel problem leverages both stochastic gradient and Hessian-vector-product (HVP) oracles to find an $\epsilon$-stationary point of $\varphi({\bm{x}})$. Subsequently, Ji et al. (2021) proposed stocBiO by incorporating multiple enhanced designs in BSA to improve the complexity. Yang et al. (2023b) proposes FdeHBO that uses finite-differences to estimate HVP vectors. However, all these methods require the stochastic Hessian assumption (5) on the lower-level function, which is stronger than the standard SGD assumption.
Kwon et al. (2023) proposed the first fully first-order method F2SA that works under standard SGD assumptions on both $f$ and $g$ (Assumption 2.1). The main idea is to solve the following penalty problem (Liu et al., 2022; 2023; Shen and Chen, 2023; Shen et al., 2025b; Lu and Mei, 2024; 2026):
@@
where $\lambda$ is taken to be sufficiently large such that $\lambda=\Omega(\epsilon^{-1})$. If we interpret $\lambda$ as the Lagrangian multiplier, then Problem (3) can be viewed as the Lagrangian function of the constrained optimization $\min_{{\bm{x}}\in{\mathbb{R}}^{d_{x}},{\bm{y}}\in{\mathbb{R}}^{d_{y}}}f({\bm{x}},{\bm{y}}),{\rm s.t.}~g({\bm{x}},{\bm{y}})\leq g({\bm{x}},{\bm{y}}^{*}({\bm{x}}))$. Thanks to Danskin’s theorem, the gradient of the Problem (3) only involves gradient information. Therefore, F2SA does not require the stochastic Hessian assumptions (5). More importantly, by directly leveraging gradient oracles instead of more expensive HVP oracles, the F2SA is more efficient in practice (Shen et al., 2025a; Xiao and Chen, 2025; Jiang et al., 2025) and is also the only method that can be scaled to 3232B sized large language model (LLM) training (Pan et al., 2024).
Kwon et al. (2023) proved that the F2SA method finds an $\epsilon$-stationary point of $\varphi({\bm{x}})$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls in the deterministic case and $\tilde{\mathcal{O}}(\epsilon^{-7})$ stochastic first-order oracle (SFO) calls in the stochastic case. Recently, Chen et al. (2025b) showed the two-time-scale stepsize strategy improves the upper complexity bound of F2SA method to $\tilde{\mathcal{O}}(\epsilon^{-2})$ in the deterministic case, which is optimal up to logarithmic factors. However, the direct extension of their method in the stochastic case leads to the $\tilde{\mathcal{O}}(\epsilon^{-6})$ SFO complexity (Chen et al., 2025b; Kwon et al., 2024a) , which still has a significant gap between the $\Omega(\epsilon^{-4})$ lower bound for SGD (Arjevani et al., 2023). It remains open whether optimal rates for stochastic bilevel problems can be achieved for fully first-order methods.
@@
Compared with Eq. (3), this new penalty problem perturbs the lower-level variables $\bm{y}$} and $\bm{z}$} in the opposite direction to better cancel out the approximation errors to Problem (1). The connection between bilevel optimization and finite difference approximations was recently established by Chayti and Jaggi (2024) in the context of meta-learning, but their results were limited to symmetric approximations. We extend their findings beyond meta-learning and to general finite difference approximations, addressing their conjecture about broader applicability. In this work, we show that F2SA-2 provably improves the SFO complexity of F2SA from $\tilde{\mathcal{O}}(\epsilon^{-6})$ to $\tilde{\mathcal{O}}(\epsilon^{-5})$ for second-order smooth problems. Moreover, our idea is generalizable for any $p$th-order smooth problems. It is known in numerical analysis there exists the $p$th-order central difference that uses $p$ points to construct an estimator to the derivative of a unitary function with $p$th-order error guarantee, as recalled in Lemma 3.1. Motivated by this fact, we propose the F2SA-$p$ algorithm and show that it enjoys the improved $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ SFO complexity, as formally stated in Theorem 3.1.
Moreover, as formally stated in Theorem 4.1, it is easy to extend the $\Omega(\epsilon^{-4})$ lower bound for SGD (Arjevani et al., 2023) from single-level optimization to bilevel optimization using a fully separable construction that automatically satisfies all our additional smoothness conditions in Definition 2.2. It shows that F2SA-$p$ is optimal up to logarithmic factors when $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$ (see Remark 3.4). We summarize our main results in Table 1 and discuss open problems in the following.
@@
Our upper bounds improve known results for high-order smooth problems, but our result still has a gap between the $\Omega(\epsilon^{-4})$ lower bound for $p={\mathcal{O}}(\log\epsilon^{-1}/\log\log\epsilon^{-1})$. Recently, Kwon et al. (2024a) obtained some preliminary results towards closing this gap for $p=1$, where they showed an $\Omega(\epsilon^{-6})$ lower bound holds under a more adversarial oracle. But it is still open whether their lower bounds can be extended to standard stochastic oracles. Another open problem is the tightness of the condition number dependency, for which the current upper and lower bounds have a gap of $\Omega(\kappa^{9})$ as demonstrated in Table 1. Two recent concurrent works (Ji, 2025; Chen and Zhang, 2025) proposed tighter lower bounds for $p=1$: Ji (2025) showed a lower bound of $\Omega(\kappa^{5/2}\epsilon^{-4})$ and Chen and Zhang (2025) showed that of $\Omega(\kappa^{4}\epsilon^{-4})$. However, it is still open to fully close the gap in condition number dependency for both $p=1$ and $p\geq 2$.
@@
F2SA 1st-order (Kwon et al., 2023) $\tilde{\mathcal{O}}({\rm poly}(\kappa)\epsilon^{-7})$
F2SA 1st-order (Kwon et al., 2024a) $\tilde{\mathcal{O}}({\rm poly}(\kappa)\epsilon^{-6})$
F2SA 1st-order (Chen et al., 2025b) $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$
[…]
@@ 2 Preliminaries
The goal of bilevel optimization is to minimize the hyper-objective $\varphi({\bm{x}})$, which is in general nonconvex. Since finding a global minimizer of a general nonconvex function requires exponential complexity in the worst case (Nemirovskij and Yudin, 1983, § 1.6), we follow the literature (Carmon et al., 2020; 2021) to consider the task of finding an approximate stationary point.
@@
Let $\varphi:{\mathbb{R}}^{d_{x}}\rightarrow{\mathbb{R}}$ be the hyper-objective defined in Eq. (1). We say a random variable $\hat{\bm{x}}\in{\mathbb{R}}^{d_{x}}$ is an $\epsilon$-hyper-stationary point if $\mathbb{E}\|\nabla\varphi(\hat{\bm{x}})\|\leq\epsilon$.
Next, we introduce the assumptions used in this paper, which ensure the tractability of the above hyper-stationarity. Compared to (Kwon et al., 2023; Chen et al., 2025b), we additionally assume the high-order smoothness in lower-level variable $\bm{y}$} to achieve further acceleration.
@@ 2.1 Problem Setup
First of all, we follow the standard assumptions on SGD (Arjevani et al., 2023) to assume that the stochastic gradient estimators satisfy the following assumption.
@@
Third, we assume the lower-level function $g({\bm{x}},{\bm{y}})$ is strongly convex in $\bm{y}$}. It guarantees the uniqueness of ${\bm{y}}^{*}({\bm{x}})$ and the tractability of the bilevel problem. Although not all the problems in applications satisfy the lower-level strong convexity assumption, it is impossible to derive dimension-free upper bounds when the lower-level problem is only convex (Chen et al., 2024, Theorem 3.2). Hence, we follow most existing works to consider strongly convex lower-level problems.
@@
Fourth, we require the following smoothness assumptions following (Ghadimi and Wang, 2018). According to Eq. (2), these conditions are necessary and sufficient to guarantee the Lipschitz continuity of $\nabla\varphi({\bm{x}})$, which further ensure the tractability of an approximate stationary point of the nonconvex hyper-objective $\varphi({\bm{x}})$ (Zhang et al., 2020; Kornowski and Shamir, 2022).
@@
We refer to the problem class that jointly satisfies all the above Assumption 2.1, 2.2, 2.3 and 2.4 as first-order smooth bilevel problems, for which (Kwon et al., 2024a; Chen et al., 2025b) showed the F2SA method achieves the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper complexity bound. In this work, we show an improved bound under the following additional higher-order smoothness assumption on lower-level variable $\bm{y}$}.
Assumption 2.5 (High order smoothness in $\bm{y}$}).
@@
All our above assumptions align with (Chen et al., 2025b) except for the additional Assumption 2.5. A classic example of a highly smooth function is the softmax function (Garg et al., 2021, Lemma 2(3)). Therefore, many hyper-parameter tuning problems for logistic regression are provably highly smooth, such that our theory can be applied. We give two examples from (Pedregosa, 2016): the first one aims to learn the optimal weights of each sample in a corrupted training set, and the second one aims to learn the optimal regularizer of each parameter.
Example 2.1 (Data hyper-cleaning).
Let ${\bm{x}}\in{\mathbb{R}}^{n}$ parameterize the per-sample weight of a training set with $n$ samples via $\sigma(x_{i})=\exp(x_{i})/\sum_{i=1}^{n}\exp(x_{i})$ and ${\bm{y}}\in{\mathbb{R}}^{d}$ be the parameters of a linear model. Let $\ell_{\rm val}$ be the logistic loss of the linear model on the validation set and $\ell_{\rm tr}^{i}$ be the logistic loss on the training sample $i$. The problem aims to solve
$\min_{{\bm{x}}\in{\mathbb{R}}^{n}}\ell_{\rm val}({\bm{y}}),\quad{\rm s.t.}\quad{\bm{y}}\in\arg\min_{{\bm{y}}\in{\mathbb{R}}^{d}}\sum_{i=1}^{n}\sigma(x_{i})\ell_{\rm tr}^{i}({\bm{y}}).$
Example 2.2 (Learn-to-regularize).
Let ${\bm{x}}\in{\mathbb{R}}^{d}$ parameterize the regularization matrix via ${\bm{W}}_{\bm{x}}={\rm diag}(\exp({\bm{x}}))$, and ${\bm{y}}\in{\mathbb{R}}^{d}$ be the parameters of a linear model. Let $\ell_{\rm val}$ and $\ell_{\rm tr}$ be the logistic loss of the linear model on the validation set and training set, respectively. The problem aims to solve
$\min_{{\bm{x}}\in{\mathbb{R}}^{d}}\ell_{\rm val}({\bm{y}}),\quad{\rm s.t.}\quad{\bm{y}}\in\arg\min_{{\bm{y}}\in{\mathbb{R}}^{d}}\ell_{\rm tr}({\bm{y}})+{\bm{y}}^{\top}{\bm{W}}_{{\bm{x}}}{\bm{y}}.$
@@ 2.2 Comparison to Previous Works
Before we show our improved upper bound, we first give a detailed discussion on other assumptions made in related works.
@@
Ghadimi and Wang (2018); Ji et al. (2021) assumes the access to a stochastic Hessian estimator ${\bm{H}}({\bm{x}},{\bm{y}})$ such that
@@
According to (Arjevani et al., 2020, Observation 1 and 2), such an assumption is stronger than standard SGD assumptions and equivalent to the mean-squared-smoothness assumption (6) on the lower-level gradient estimator $G$ under the mild condition of $\nabla G({\bm{x}},{\bm{y}})={\bm{H}}({\bm{x}},{\bm{y}})$. Under this assumption, in conjunction with Assumption 2.2, 2.3, and 2.4, Ghadimi and Wang (2018) proposed the BSA method that can find an $\epsilon$ stationary point of $\varphi({\bm{x}})$ with $\tilde{\mathcal{O}}(\epsilon^{-6})$ SFO complexity and $\tilde{\mathcal{O}}(\epsilon^{-4})$ stochastic HVP complexity. Later, Ji et al. (2021) further improved the SFO complexity term to $\tilde{\mathcal{O}}(\epsilon^{-4})$. Compared to them, we consider the setting where the algorithms only have access to stochastic gradient estimators, and make no assumptions on the stochastic Hessians.
@@
Besides Assumption 2.1, 2.2, 2.3, 2.4 and the stochastic Hessian assumption (5), Khanduri et al. (2021); Yang et al. (2021; 2023b) further assumes that the stochastic estimators to gradients and Hessians are mean-squared smooth:
@@
Under this additional assumption, they proposed faster stochastic methods with upper complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-3})$ via variance reduction (Fang et al., 2018; Cutkosky and Orabona, 2019). In this paper, we only consider the setting without mean-squared smoothness assumptions and study a different acceleration mechanism from variance reduction.
@@
Huang et al. (2025) introduced a second-order smoothness assumption similar to but stronger than Assumption 2.5 when $p=2$. Specifically, they assumed the second-order smoothness jointly in $({\bm{x}},{\bm{y}})$ instead of $\bm{y}$} only:
@@
The jointly second-order smoothness (7) ensures that the hyper-objective $\varphi({\bm{x}})$ has Lipschitz continuous Hessians, which further allows the application of known techniques in minimizing second-order smooth objectives. Huang et al. (2025) applied the technique from (Jin et al., 2017; 2021; Xu et al., 2018; Allen-Zhu and Li, 2018) to show that an HVP-based method can find a second-order stationary point in $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity under the deterministic setting, and in $\tilde{\mathcal{O}}(\epsilon^{-4})$ under the stochastic Hessian assumption (5). Yang et al. (2023a) applied the technique from (Li and Lin, 2023) to accelerate the complexity HVP-based method to $\tilde{\mathcal{O}}({\epsilon}^{-1.75})$ in the deterministic setting. Chen et al. (2025b) also proposed a fully first-order method to achieve the same $\tilde{\mathcal{O}}({\epsilon}^{-1.75})$ complexity. Compared to these works, our work demonstrates a unique acceleration mechanism in stochastic bilevel optimization that only comes from the high-order smoothness in $\bm{y}$}.